# Privacy-Preserving in Connected and Autonomous Vehicles Through Vision to Text Transformation

### v1

@@ 2 RELATED WORK
3 PROBLEM STATEMENT AND PROPOSED SOLUTION
@@ 4 RESULTS AND DISCUSSIONS
Figure 3: Image Reconstruction Metrics.
Figure 4: Detailed-Oriented Metrics.
The experimental framework is implemented in Python 3.11, where input images are processed to generate descriptive text outputs. To evaluate the generalizability of the proposed model, two datasets are employed: Drive&Act [20], which contains approximately 9.6 million images, and a public traffic-scene dataset from Kaggle [21]. These datasets cover diverse environments to assess the model’s performance under varying conditions.
@@ 4.1 Privacy Preservation Metrics
To check how well privacy is protected, we use a test called the Adversarial Reconstruction Test. In this test, pictures are recreated from the text that was made and then checked against the original pictures. The following metrics are used: (1) Mean Squared Error (MSE): calculates how much each pixel in the reconstructed image differs from the corresponding pixel in the original image. Although MSE doesn’t measure privacy directly, a bigger MSE value means the recreated image has more differences from the original, implying it has less private visual details. (2) Structural Similarity Index Measure (SSIM): evaluates how similar two images appear to the human eye by looking at brightness, contrast, and the arrangement of features. When the SSIM score is low (e.g., less than 0.5), it means the image doesn’t look much like the original one. In our findings (Figure  3), the model effectively maintains visual privacy, as indicated by the fact that the MSE reaches $10^{4}$ and SSIM falls well below 0.5, showing the model’s ability to preserve visual privacy. This assessment structure determines the difficulty of reorganizing the original visual data from the output text, thus reflecting the effectiveness of the model in protection of privacy.
@@ 4.2 Text Quality and Semantic Richness Metrics
The information and linguistic quality of the generated text is assessed by using the following literal and meaning metrics: (1) Named Entities and Modifiers: evaluates whether the model correctly captures specific entities (e.g., people, places, objects) and uses rich descriptive modifiers. As illustrated in Figure  4, the baseline model often leaves such elements, while the proposed model consistently considers more. (2) Word Count and Unique Word Count: reflect the length and literal variety of the output. As shown in Figure  5, the proposed model produces a longer and more detailed description than the baseline, and quantitatively indicates better expression. (3) Detailed Density: determines the amount of specific and descriptive information relative to the overall length of the text. It captures semantic prosperity by identifying constructions such as subordinate sections and detailed comments. Figure  6 highlights the proposed model produces intensive and more informative details, contributing to more accurate visual understanding. (4) Semantic Similarity Score: compares two texts generated from the same image to assess the semantic similarity. The score of 0.7371 was obtained, confirming that the proposed model has high meaning alignment with baseline offering richer descriptive text. Figure 7 illustrates the semantic similarity between the text generated by the baseline model and the proposed model. Initially, the proposed model produces descriptions that are closely aligned with those of the baseline. However, as training progresses, the model receives updates and improvements. After 500 iterations, it reaches a stability, with the baseline model’s output remaining approximately 73% semantically similar to the text generated by the proposed model. This demonstrates that the proposed model is capable of enriching the generated text with additional semantic content.
Although these matrices do not measure direct privacy, they are important to assess the interpretation, utility and relevance of the text generated for downstream functions.
@@
Figure 7: Semantic Similarity Score, which indicates how similar the text generated by the baseline method is to that of the proposed one.

### v2

@@ 2 RELATED WORK
3 PROPOSED SOLUTION
@@ 4 RESULTS AND DISCUSSIONS
The experimental framework is implemented in Python 3.11, where input images are processed to generate descriptive text outputs. To evaluate the generalizability of the proposed model, two datasets are employed: Drive&Act [20], which contains approximately 9.6 million images, and a public traffic-scene dataset from Kaggle [21]. These datasets cover diverse environments to assess the model’s performance under varying conditions. Initially the model will be compared with the results obtained from a PPO-based RL model then it will be compared to the model introduced in [22]. Therefore, we also consider CFP-FP and AgeDB-30 datasets to apply .
@@
Table 2: Privacy Metrics Evaluation on the CFP-FP Dataset
Method SSIM $\downarrow$ PSNR $\downarrow$ MSE $\downarrow$ SRRA (%) $\downarrow$
AdvFace 0.89 23.54 314.7 13.01
PPO-based RL 0.86 23.02 333.7 12.5
Feedback-based (Ours) 0.85 22.18 365.2 11.25
Table 3: Privacy Metrics Evaluation on the AgeDB-30 Dataset
Method SSIM $\downarrow$ PSNR $\downarrow$ MSE $\downarrow$ SRRA (%) $\downarrow$
AdvFace 0.87 22.91 347.8 14.12
PPO-based 0.86 21.98 355.5 13.03
Feedback-based (Ours) 0.83 21.88 389.1 12.63
Figure 3: NER and Modifiers Metrics.
@@ 4.1 Privacy Preservation Metrics
To check how well privacy is protected, we use a test called the Adversarial Reconstruction Test. In this test, pictures are reconstructed and then compared against the original pictures. The following metrics are used where Table 2 and 3 presents results for the first four metrics, metrics in 5 are shown in Figure 3:
  1. 1.
    SSIM evaluates how similar two images are.
  2. 2.
    PSNR metric evaluates the quality of a compressed or reconstructed image by calculating the ratio between the maximum possible signal power and the power of the corrupting noise.
  3. 3.
    MSE calculates how much each pixel in the reconstructed image differs from the corresponding pixel in the original image.
  4. 4.
    SRRA to measure the performance of the protected features.
  5. 5.
    Named Entities and Modifiers evaluates whether the model correctly captures specific entities (e.g., people, places, objects) and uses rich descriptive modifiers. As illustrated in Figure  3, the baseline model often leaves such elements, while the proposed model consistently considers more.
@@ 4.2 Text Quality and Semantic Richness Metrics
The information and linguistic quality of the generated text is assessed by using the following literal and meaning metrics:
  • Word Count and Unique Word Count: reflect the length and literal variety of the output. As shown in Figure  4, the proposed model produces a longer and more detailed description than the baseline, and quantitatively indicates better expression.
  • Detailed Density: determines the amount of specific and descriptive information relative to the overall length of the text. It captures semantic prosperity by identifying constructions such as subordinate sections and detailed comments. Figure  5 highlights the proposed model produces intensive and more informative details, contributing to more accurate visual understanding.
  • Semantic Similarity Score: compares two texts generated from the same image to assess the semantic similarity. The score of 0.7371 was obtained, confirming that the proposed model has high meaning alignment with baseline offering richer descriptive text.
    Figure 6 illustrates the semantic similarity between the text generated by the baseline model and the proposed model. Initially, the proposed model produces descriptions that are closely aligned with those of the baseline. However, as training progresses, the model receives updates and improvements. After 500 iterations, it reaches a stability, with the baseline model’s output remaining approximately 73% semantically similar to the text generated by the proposed model. This demonstrates that the proposed model is capable of enriching the generated text with additional semantic content.
@@
Figure 6: Semantic Similarity Score, which indicates how similar the text generated by the baseline method is similar to that of the proposed one.